The probability that a person owns a microwave oven is 0.75. The probability that a person owns a compact disk player is 0.25. The probability that a person owns both a microwave and a compact disk player is 0.16. Find the probability that a person owns either a microwave oven or a compact disk player but not both.
Answer #1

P(microwave or compact) = P(microwave) + P(compact) - P(microwave and compact)

= 0.75 + 0.25 - 0.16

= 0.84
Answer #2

P = 0.75 + 0.25 - 0.16 P = 0.84

answered by: anonymous